1) (S)(-) mandelic acid has a specific rotation of -158.0 degrees, If the rotation you observed in your polarimeter in a 2.5 dm cell was -150.1, what is the concentration of the solution in g ml1?
Answer #1

[X] = X/Cl

[x] = specific rotation = -158.0 degrees

x = optical rotation = -150.1 degrees

C = concentration in g/ml = ?

l = length of the sample tube = 2.5 dm

-158 = -150.1/(2.5*C)

C = 0.38 g/ml